Grade 8Physical Chemistry

calculate the kinetic energies of 3gms of hydrogen and 4gms of oxygen at a given temperature

mgayathri
10 Years agoGrade 8

1 Answer

Vikas TU
10 Years ago
Since,  H2 and O2 are ideal gases,
the kinetic aEnergies for gases is given as:
3/2nRT
T is the specific temperature..
 
Just calculate the moles for the gases.
  1. For H2 moles = 3/2 =  1.5
  2. For O2 moles = 4/32 = 1/8
Now U can Calculate.